# Trace Reconstruction of First-Order Reed-Muller Codewords Using Run Statistics

### v1

@@ Abstract
In this paper, we derive an expression for the expected number of runs in a trace of a binary sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$ obtained by passing $x$ through a deletion channel that independently deletes each bit with probability $q$. We use this expression to show that if $x$ is a codeword of a first-order Reed-Muller code, and the deletion probability $q$ is $1/2$, then $x$ can be reconstructed, with high probability, from $\widetilde{O}(n)$ many of its traces.
@@ I Introduction
In the trace reconstruction problem, a sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$ is to be reconstructed exactly from multiple independent traces obtained by repeatedly passing $x$ through an i.i.d. deletion channel that independently deletes each bit with probability $q$. The goal is to determine the trace complexity of reconstruction, i.e., the minimum number of traces required to reconstruct $x$ with high probability (w.h.p.) and also to find an efficient reconstruction algorithm. The problem has its origins in the work of Levenshtein [1, 2], but in the form described above, it was first introduced by Batu et al. [3].
@@
To the best of our knowledge, all prior works that derive upper bounds on the trace complexity of worst-case sequence reconstruction rely on aggregating in some way the statistics of individual bits in a trace in order to differentiate between candidate sequences. In our work, we introduce a new statistic — the number of runs in a trace — and demonstrate how it can be used to differentiate between the codewords of $\mathrm{RM}(m,1)$. By doing so, we obtain an $\widetilde{O}(n)$ upper bound on the (worst-case) trace complexity of sequences from $\mathrm{RM}(m,1)$ for a deletion probability of $q=\frac{1}{2}$. While runs of $\mathsf{0}$s and $\mathsf{1}$s have previously been used in reconstruction algorithms [3], [10], [11], they were invariably used as an aid in aligning the bits of a trace to the bits of the original sequence $x$. In contrast, our use of the number of runs in a trace is purely statistical.
This paper has two main contributions. The first is Theorem 1, presented in Section III, which gives an expression for the expected number of runs in a (random) trace of a given sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$. The other is Theorem 2 (Section IV), which states that the trace complexity of reconstructing w.h.p. an arbitrary codeword from $\mathrm{RM}(m,1)$ is $\widetilde{O}(n)$ when the deletion probability $q$ is $\frac{1}{2}$.
@@ II Notation and Preliminaries
Let $x$ be a binary sequence of length $n$. The support of $x$ is the set $\mathcal{S}_{x}=\{i:x_{i}=\mathsf{1}\}$. We will use $\overline{\mathcal{S}}_{x}$ to denote the complementary set $[n]\setminus\mathcal{S}_{x}=\{i:x_{i}=\mathsf{0}\}$. The Hamming weight of $x$ is $w_{H}(x)=|\mathcal{S}_{x}|$. A run of $\mathsf{0}$s of length $l$ in $x$ is a subsequence of the form $(x_{i},x_{i+1},\cdots,x_{i+l-1})$, $11\leq i\leq n-l+$ 1, with $x_{i-1}=\mathsf{1},x_{i}=x_{i+1}=\cdots=x_{i+l-1}=\mathsf{0},x_{i+l}=\mathsf{1}$, where to handle edge cases, we set $x_{0}=x_{n+1}=\mathsf{1}$. A run of $\mathsf{1}$s is defined analogously.
@@ II-A Binary First-Order Reed-Muller Codes
In this section, we record some standard facts about binary first-order Reed-Muller (RM) codes that will be used later in the paper; for more information on RM codes, see, e.g., [14]. Let $\mathbb{F}_{2}$ denote the binary field. For a fixed integer $m\geq 1$, let $\mathsf{x}_{1},\mathsf{x}_{2},\ldots,\mathsf{x}_{m}$ be $m$ Boolean variables. For a polynomial $f\in\mathbb{F}_{2}[\mathsf{x}_{1},\mathsf{x}_{2},\ldots,\mathsf{x}_{m}]$, and a binary $m$-tuple $z=(z_{1},z_{2},\ldots,z_{m})\in\mathbb{F}_{2}^{m}$, the evaluation of $f$ at $z$ is $\text{Eval}_{z}(f)=f(z_{1},z_{2},\cdots,z_{m})$, all arithmetic taking place in $\mathbb{F}_{2}$. Then, we define the evaluation vector $\text{Eval}(f):=(\text{Eval}_{z}(f):z\in\mathbb{F}_{2}^{m})$, with the evaluation points $z\in\mathbb{F}_{2}^{m}$ appearing in the standard lexicographic order. The binary first-order RM code $\mathrm{RM}(m,1)$ is the set of all evaluation vectors $\text{Eval}(f)$ of polynomials $f$ of the form $u_{0}+u_{1}\mathsf{x}_{1}+u_{2}\mathsf{x}_{2}+\cdots+u_{m}\mathsf{x}_{m}$, with $u_{i}\in\mathbb{F}_{2}$ for all $i$. The code has blocklength $n=2^{m}$, dimension $m+1$, and minimum distance $n/2=2^{m-1}$. In fact, the code consists of the all-zeros and all-ones vectors, $n-1$ vectors of Hamming weight $n/2$ with first coordinate $\mathsf{0}$, and $n-1$ vectors of Hamming weight $n/2$ with first coordinate $\mathsf{1}$. In particular, since the all-ones vector is a codeword, $x\in\mathrm{RM}(m,1)$ implies that $\overline{x}\in\mathrm{RM}(m,1)$.
@@ II-B Chernoff Bounds
  • for $\delta>0$,  $\Pr(X\geq(1+\delta)\mu)\leq e^{-\mu\delta^{2}/(2+\delta)}$; in particular, for $10<\delta<$ 1, $\Pr(X\geq(1+\delta)\mu)\leq e^{-\mu\delta^{2}/3}$
@@ IV Trace Reconstruction for $\mathrm{RM}(m,1)$
From here on, we consider only the case of deletion probability $q=\frac{1}{2}$. Our main result is that, using the average number of runs in a trace as a statistic, one can reconstruct, with high probability (w.h.p.), any $x\in\mathrm{RM}(m,1)$ using only $\widetilde{O}(n)$ of its traces, where $n=2^{m}$ is the blocklength of the code.
@@
For deletion probability $q=\frac{1}{2}$, any codeword $x\in\mathrm{RM}(m,1)$ can be correctly reconstructed from $\widetilde{O}(n)$ traces, with probability $\geq 1-1/\mathrm{poly}(n)$.
We first describe the proof strategy. Recall that $\mathrm{RM}(m,1)$ consists of $2n=2^{m+1}$ codewords, of which half begin with a $\mathsf{0}$ and the other half begin with a $\mathsf{1}$. By examining the first bit of $O(\log n)$ traces, the first bit of $x$ can be correctly determined w.h.p. — see Claim 1 below. It then remains to distinguish between the $n$ codewords of having $x_{1}$ as the first bit. This, as we will show, can be achieved w.h.p. using run statistics obtained from $O(n\log n)$ traces. What makes this possible is the following property of the $\mathrm{RM}(m,1)$ code, which may be of independent interest.
@@ IV-A Proof of Theorem 2
We will prove that if the algorithm takes $k=O(n\log n)$ random traces as input, and we set $\ell=O(\log n)$ in Step 2, then with probability at least $1-1/\mathrm{poly}(n)$, the codeword output by the algorithm is equal to the $x$ from which the $k$ traces were obtained. In what follows, $\mathbb{P}_{x}$ denotes the probability law of the traces $T_{x}^{1},T_{x}^{2},\ldots,T_{x}^{k}$, and $Z$ is the random variable representing the output of the reconstruction algorithm. Reconstruction Error is the event $Z\neq x$, and $\mathcal{E}_{j}$, for $j=1,2$, is the event “Error occurs in Step $j$”. Note that
@@
Suppose that the first bit of $x\in\mathrm{RM}(m,1)$ is $\mathsf{1}$. If the second bit is also $\mathsf{1}$, then $\mathbb{E}[A_{i}]\geq 1-\mathbb{P}_{x}($the first two bits of $x$ are deleted)=3/4)=3/4) = 3 / 4. If the second bit of $x$ is $\mathsf{0}$, then the next two bits are either $\mathsf{1}\mathsf{0}$ or $\mathsf{0}\mathsf{1}$ (see Section II-A). In this case, $\mathbb{E}[A_{i}]\geq\mathbb{P}_{x}($one of the two $\mathsf{1}$s in the first four bits of $x$ is the first bit of trace $T_{x}^{i})\geq(1/2+1/16)=9/16$. Thus, when the first bit of $x$ is $\mathsf{1}$, then $\mathbb{E}[A]\geq\frac{9}{16}\ell$. Via the Chernoff bound, we obtain $\mathbb{P}_{x}(\mathcal{E}_{1})=\mathbb{P}_{x}(A\leq\frac{1}{2}\ell)\leq e^{-\frac{\ell}{288}}$.
@@
If $k=O(n\log n)$ traces are given as input to the reconstruction algorithm, then for any $x\in\mathrm{RM}(m,1)$, we have $\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})\leq 2\cdot n^{-c_{2}}$ for some constant $c_{2}>0$.
@@
Since $\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})\leq\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})$, it suffices to bound the latter probability. Let $\mathbb{E}[R_{x}]=r_{x}\leq n$. From Proposition 1, for any other $y\in\mathrm{RM}(m,1)$ having the same first bit as $x$, we have $|r_{x}-\mathbb{E}[R_{y}]|\geq\Delta:=0.028$. So, if the reconstruction algorithm does not make an error in Step 1 (i.e., given $\mathcal{E}_{1}^{c}$), then at the end of Step 2, it will output $Z=x$ if $r_{x}-\Delta/2<\overline{R}<r_{x}+\Delta/2$. Consequently, given $\mathcal{E}_{1}^{c}$, if error event $\mathcal{E}_{2}$ occurs, then either $\overline{R}\leq r_{x}-\Delta/2$ of $\overline{R}\geq r_{x}+\Delta/2$. Thus,
$\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})\leq\mathbb{P}_{x}(\overline{R}\leq r_{x}-\Delta/2)+\mathbb{P}_{x}(\overline{R}\geq r_{x}+\Delta/2).$
The two probabilities on the right-hand side will be bounded using the Chernoff bound.
Note that $k\overline{R}=\sum_{i=1}^{k}R_{x}^{i}=\sum_{i=1}^{k}(R_{x,\mathsf{0}}^{i}+R_{x,\mathsf{1}}^{i})$. Since $R_{x,\mathsf{0}}$ and $R_{x,\mathsf{1}}$ are expressible as sums of indicator random variables (as in the proof of Theorem 1), the Chernoff bounds in Section II-B can be applied to obtain
$\mathbb{P}_{x}\bigl{(}k\overline{R}\leq k(r_{x}-\Delta/2)\bigr{)}$ $\leq e^{-\frac{k\cdot\Delta^{2}}{8r_{x}}}$
[…]
Hence, $\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})\leq 2\,e^{-\frac{k\cdot\Delta^{2}}{12r_{x}}}$. Taking $k=c_{3}n\log n$ (for suitable $c_{3}>0$) we obtain $\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})\leq 2\cdot e^{-\frac{c_{3}n\log n\cdot\Delta^{2}}{12r_{x}}}\leq 2\cdot e^{-\frac{c_{3}\Delta^{2}}{\mathsf{1}2}\log n}=2\cdot n^{-c_{2}}$. ∎
@@ IV-B Proof of Proposition 1
$\beta_{x}$ $=\ \sum_{i,j\in\overline{\mathcal{S}}_{x}:i<j}\biggl{(}\frac{1}{2}\biggr{)}^{n-(j-i)}+\sum_{i,j\in{\mathcal{S}}_{x}:i<j}\biggl{(}\frac{1}{2}\biggr{)}^{n-(j-i)}$
$\gamma_{x}$ $=\ \sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}\biggl{(}\frac{1}{2}\biggr{)}^{n-(j-i)}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}\biggl{(}\frac{1}{2}\biggr{)}^{n-(j-i)}$
$\delta_{x}$ $=\ \sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}\biggl{(}\frac{1}{2}\biggr{)}^{j-i}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}\biggl{(}\frac{1}{2}\biggr{)}^{j-i}.$
@@
[…]

### v3

@@ Abstract
In this paper, we derive an expression for the expected number of runs in a trace of a binary sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$ obtained by passing $x$ through a deletion channel that independently deletes each bit with probability $q$. We use this expression to show that if $x$ is a codeword of a first-order Reed-Muller code, and the deletion probability $q$ is $1/2$, then $x$ can be reconstructed, with high probability, from $\widetilde{O}(n^{2})$ many of its traces.
@@ I Introduction
In the trace reconstruction problem, a sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$ is to be reconstructed exactly from multiple independent traces obtained by repeatedly passing $x$ through an i.i.d. deletion channel that independently deletes each bit with probability $q$. The goal is to determine the trace complexity of reconstruction, i.e., the minimum number of traces required to reconstruct $x$ with high probability (w.h.p.) and also to find an efficient reconstruction algorithm. The problem has its origins in the work of Levenshtein [Lev97, Lev01], but in the form described above, it was first introduced by Batu et al. [BKKM04].
@@
Our approach is in the spirit of that of Chase [Chase20], which uses a statistic based on the appearances of a specific subsequence in the observed traces to distinguish between a pair of sequences $x$ and $y$. We use the simpler notion of runs (i.e., contiguous subsequences of like symbols) to craft our statistic. Specifically, we demonstrate that the number of runs of $\mathsf{0}$s and $\mathsf{1}$s in the observed traces can be used to differentiate between the codewords of $\mathrm{RM}(m,1)$. Doing so, we obtain an $\widetilde{O}(n^{2})$ upper bound on the (worst-case) trace complexity of sequences from $\mathrm{RM}(m,1)$ for a deletion probability of $q=\frac{1}{2}$. While runs of $\mathsf{0}$s and $\mathsf{1}$s have previously been used in reconstruction algorithms [BKKM04], [CGMR20], [KMMP21], they were invariably used as an aid in aligning the bits of a trace to the bits of the original sequence $x$. In contrast, our use of the number of runs in a trace is purely statistical.
This paper has two main contributions. The first is Theorem 1, presented in Section III, which gives an expression for the expected number of runs in a (random) trace of a given sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$. The other is Theorem 2 (Section IV), which states that the trace complexity of reconstructing w.h.p. an arbitrary codeword from $\mathrm{RM}(m,1)$ is $\widetilde{O}(n^{2})$ when the deletion probability $q$ is $\frac{1}{2}$.
@@ II Notation and Preliminaries
Let $x$ be a binary sequence of length $n$. The support of $x$ is the set $\mathcal{S}_{x}=\{i:x_{i}=\mathsf{1}\}$. We will use $\overline{\mathcal{S}}_{x}$ to denote the complementary set $n]\setminus\mathcal{S}_{x}=\{i:x_{i}=\mathsf{0}\}$. The Hamming weight of $x$ is $w_{H}(x)=|\mathcal{S}_{x}|$. A run of $\mathsf{0}$s of length $l$ in $x$ is a subsequence of the form $(x_{i},x_{i+1},\cdots,x_{i+l-1})$, $1\leq i\leq n-l+1$, with $x_{i-1}=\mathsf{1},x_{i}=x_{i+1}=\cdots=x_{i+l-1}=\mathsf{0},x_{i+l}=\mathsf{1}$, where to handle edge cases, we set $x_{0}=x_{n+1}=\mathsf{1}$. A run of $\mathsf{1}$s is defined analogously.
@@ II-A Binary First-Order Reed-Muller Codes
In this section, we record some standard facts about binary first-order Reed-Muller (RM) codes that will be used later in the paper; for more information on RM codes, see, e.g., [ASY21]. Let $\mathbb{F}_{2}$ denote the binary field. For a fixed integer $m\geq 1$, let $\mathsf{x}_{1},\mathsf{x}_{2},\ldots,\mathsf{x}_{m}$ be $m$ Boolean variables. For a polynomial $f\in\mathbb{F}_{2}[\mathsf{x}_{1},\mathsf{x}_{2},\ldots,\mathsf{x}_{m}]$, and a binary $m$-tuple $z=(z_{1},z_{2},\ldots,z_{m})\in\mathbb{F}_{2}^{m}$, the evaluation of $f$ at $z$ is $\text{Eval}_{z}(f)=f(z_{1},z_{2},\cdots,z_{m})$, all arithmetic taking place in $\mathbb{F}_{2}$. Then, we define the evaluation vector $\text{Eval}(f):=(\text{Eval}_{z}(f):z\in\mathbb{F}_{2}^{m})$, with the evaluation points $z\in\mathbb{F}_{2}^{m}$ appearing in the standard lexicographic order. The binary first-order RM code $\mathrm{RM}(m,1)$ is the set of all evaluation vectors $\text{Eval}(f)$ of polynomials $f$ of the form $u_{0}+u_{1}\mathsf{x}_{1}+u_{2}\mathsf{x}_{2}+\cdots+u_{m}\mathsf{x}_{m}$, with $u_{i}\in\mathbb{F}_{2}$ for all $i$. The code has blocklength $n=2^{m}$, dimension $m+1$, and minimum distance $n/2=2^{m-1}$. In fact, the code consists of the all-zeros and all-ones vectors, n1n-1 vectors of Hamming weight $n/2$ with first coordinate $\mathsf{0}$, and n1n-1 vectors of Hamming weight $n/2$ with first coordinate $\mathsf{1}$. In particular, since the all-ones vector is a codeword, $x\in\mathrm{RM}(m,1)$ implies that $\overline{x}\in\mathrm{RM}(m,1)$.
@@ II-B Chernoff Bounds
  • for $\delta>0$, $\Pr(X\geq(1+\delta)\mu)\leq e^{-\mu\delta^{2}/(2+\delta)}$; in particular, for $0<\delta<1$, $\Pr(X\geq(1+\delta)\mu)\leq e^{-\mu\delta^{2}/3}$
@@
II-C Hoeffding’s Inequality
Let $X_{1},\cdots,X_{n}$ be independent random variables such that $a_{i}\leq X_{i}\leq b_{i}$ a.s. If $X=\sum\limits_{i=1}^{n}X_{i}$, then Hoeffding’s inequality states that
$\Pr(|X-\mathbb{E}[X]|\geq t)\leq 2\exp\biggl(-\frac{2t^{2}}{\sum\limits_{i=1}^{n}(b_{i}-a_{i})^{2}}\biggr)$
@@ IV Trace Reconstruction for $\mathrm{RM}(m,1)$
From here on, we consider only the case of deletion probability $q=\frac{1}{2}$. Our main result is that, using the average number of runs in a trace as a statistic, one can reconstruct, with high probability (w.h.p.), any $x\in\mathrm{RM}(m,1)$ using only $\widetilde{O}(n^{2})$ of its traces, where $n=2^{m}$ is the blocklength of the code.
@@
For deletion probability $q=\frac{1}{2}$, any codeword $x\in\mathrm{RM}(m,1)$ can be correctly reconstructed from $\widetilde{O}(n^{2})$ traces, with probability $\geq 1-1/\mathrm{poly}(n)$.
We first describe the proof strategy. Recall that $\mathrm{RM}(m,1)$ consists of $2n=2^{m+1}$ codewords, of which half begin with a $\mathsf{0}$ and the other half begin with a $\mathsf{1}$. By examining the first bit of $O(\log n)$ traces, the first bit of $x$ can be correctly determined w.h.p. — see Claim 1 below. It then remains to distinguish between the $n$ codewords of having $x_{1}$ as the first bit. This, as we will show, can be achieved w.h.p. using run statistics obtained from $O(n^{2}\log n)$ traces. What makes this possible is the following property of the $\mathrm{RM}(m,1)$ code, which may be of independent interest.
@@ IV-A Proof of Theorem 2
We will prove that if the algorithm takes $k=O(n^{2}\log n)$ random traces as input, and we set $\ell=O(\log n)$ in Step 1, then with probability at least $1-1/\mathrm{poly}(n)$, the codeword output by the algorithm is equal to the $x$ from which the $k$ traces were obtained. In what follows, $\mathbb{P}_{x}$ denotes the probability law of the traces $T_{x}^{1},T_{x}^{2},\ldots,T_{x}^{k}$, and $Z$ is the random variable representing the output of the reconstruction algorithm. Reconstruction Error is the event $Z\neq x$, and $\mathcal{E}_{j}$, for $j=1,2$, is the event “Error occurs in Step jj”. Note that
@@
Suppose that the first bit of $x\in\mathrm{RM}(m,1)$ is $\mathsf{1}$. If the second bit is also $\mathsf{1}$, then $\mathbb{E}[A_{i}]\geq 1-\mathbb{P}_{x}($the first two bits of $x$ are deleted)=3/4)=3/4. If the second bit of $x$ is $\mathsf{0}$, then the next two bits are either $\mathsf{1}\mathsf{0}$ or $\mathsf{0}\mathsf{1}$ (see Section II-A). In this case, $\mathbb{E}[A_{i}]\geq\mathbb{P}_{x}($one of the two $\mathsf{1}$s in the first four bits of $x$ is the first bit of trace $T_{x}^{i})\geq(1/2+1/16)=9/16$. Thus, when the first bit of $x$ is $\mathsf{1}$, then $\mathbb{E}[A]\geq\frac{9}{16}\ell$. Via the Chernoff bound, we obtain $\mathbb{P}_{x}(\mathcal{E}_{1})=\mathbb{P}_{x}(A\leq\frac{1}{2}\ell)\leq e^{-\frac{\ell}{288}}$.
@@
If $k=O(n^{2}\log n)$ traces are given as input to the reconstruction algorithm, then for any $x\in\mathrm{RM}(m,1)$, we have $\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})\leq 2\cdot n^{-c_{2}}$ for some constant $c_{2}>0$.
@@
Since $\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})\leq\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})$, it suffices to bound the latter probability. Let $\mathbb{E}[R_{x}]=r_{x}\leq n$. From Proposition 1, for any other $y\in\mathrm{RM}(m,1)$ having the same first bit as $x$, we have $|r_{x}-\mathbb{E}[R_{y}]|\geq\Delta:=0.028$. So, if the reconstruction algorithm does not make an error in Step 1 (i.e., given $\mathcal{E}_{1}^{c}$), then at the end of Step 2, it will output $Z=x$ if $r_{x}-\Delta/2<\overline{R}<r_{x}+\Delta/2$. Consequently, given $\mathcal{E}_{1}^{c}$, if error event $\mathcal{E}_{2}$ occurs, then $|\overline{R}-r_{x}|\geq\Delta/2$.
Note that $k\overline{R}=\sum_{i=1}^{k}R_{x}^{i}$. Since $0\leq R_{x}^{i}\leq n$, using Hoeffding’s inequality, we obtain
$\Pr(|\overline{R}-r_{x}|\geq\Delta/2)=\Pr(|k\overline{R}-kr_{x}|\geq k\Delta/2)\leq 2\exp\biggl(-\frac{k\Delta^{2}}{2n^{2}}\biggr)$
Taking $k=c_{3}n^{2}\log n$ (for suitable $c_{3}>0$), we obtain $\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})\leq 2\cdot e^{-\frac{c_{3}\Delta^{2}}{2}\log n}=2n^{-c_{2}}$. ∎
@@ IV-B Proof of Proposition 1
$\beta_{x}\$$=\ \sum_{i,j\in\overline{\mathcal{S}}_{x}:i<j}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}+\sum_{i,j\in{\mathcal{S}}_{x}:i<j}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}$
$\gamma_{x}\$$=\ \sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}$
$\delta_{x}\$$=\ \sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{j-i}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{j-i}.$
@@
$\bigl|\mathbb{E}[R_{x}]-\mathbb{E}[R_{y}]\bigr|\ =\ \frac{1}{2}|\alpha_{x}-\alpha_{y}|\ \geq\ \frac{1}{2}\cdot 0.057\ =\ 0.0285,$